OpenStudy (anonymous):

(x^2 - 1)/(9x)/(x^2 + 2x + 1)/(3x^2)

OpenStudy (anonymous):

simplify?

OpenStudy (anonymous):

\[\frac{\frac{x^2-1}{9x}}{\frac{x^2+2x+1}{3x^2}}\]?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

alright so do you know that \[\frac{\frac{a}{b}}{\frac{c}{d}}=\frac{a}{b}*\frac{d}{c}?\]

OpenStudy (anonymous):

basically multiplyng by one in the form of the reciprocal to get one on the bottom

OpenStudy (anonymous):

yes, i simplified it to that.

OpenStudy (anonymous):

but the answer i got wasn't in the choices, so i wanted to know what i did wrong

OpenStudy (anonymous):

\[\frac{x^2-1}{9x}*\frac{3x^2}{x^2+2x+1}\]

OpenStudy (anonymous):

did you factor \[x^2-1,3x^2,x^2+2x+1\]

OpenStudy (anonymous):

ohhh. I tried making them have the same denominators first

OpenStudy (anonymous):

before i factored

OpenStudy (anonymous):

no need to make the same denominator, you're multiplying

OpenStudy (anonymous):

oh haha i forgot about that

OpenStudy (anonymous):

alright im gonna try doing this again

OpenStudy (anonymous):

alright i'm going to go eat and 'll check the answer when i get back

OpenStudy (anonymous):

k i got x(x-1)/3(x + 1)

OpenStudy (anonymous):

right